Pleasae help me solve this word problem....Two minor league baseball players got a total of 322 hits. Washington had 8 more hits than Sanchez. find the number of hits for each player.
Answer by Theo(13342)

w = number of hits washington had.
h = number of hits sanchez had.

washington had 8 more hits than sanchez.

w = h + 8

they both got a total of 322 hits.

w + h = 322

substitute h + 8 for w to get:

(h + 8) + h = 322

simplify to get:

2h + 8 = 322

subtract 8 from both sides of the equation to get 2h = 314.

divide by 2 to get h = 157.

add 8 to get w = 165.

you have:

h = 157
w = 165.

w + h = 157 + 165 = 322 (this is good).

w = h + 8 = 157 + 8 = 165 (this is also good).